# How Good Are LLMs at Out-of-Distribution Detection?

### v3

@@ Abstract
Out-of-distribution (OOD) detection plays a vital role in enhancing the reliability of machine learning (ML) models. With the rise of large language models (LLMs), even though existing research has probed OOD detection with relatively small-scale Transformers like BERT, RoBERTa and GPT-2, the stark differences in scales, pre-training objectives, and inference paradigms call into question the applicability of these findings to LLMs. This paper embarks on a pioneering empirical investigation of OOD detection in the domain of LLMs, focusing on LLaMA series ranging from 7B to 65B in size. We thoroughly evaluate commonly used OOD detectors, scrutinizing their performance in both zero-grad and fine-tuning scenarios. Notably, we alter previous discriminative in-distribution fine-tuning into generative fine-tuning, aligning the pre-training objective of LLMs with downstream tasks. Our findings unveil that a simple cosine distance OOD detector demonstrates superior efficacy, outperforming other OOD detectors. We provide an intriguing explanation for this phenomenon by highlighting the isotropic nature of the embedding spaces of LLMs, which distinctly contrasts with the anisotropic property observed in smaller BERT family models. The new insight enhances our understanding of how LLMs detect OOD data, thereby enhancing their adaptability and reliability in dynamic environments. We have released the source code at https://github.com/Awenbocc/LLM-OOD for other researchers to reproduce our results.
@@
[…]
Department of Computing, The Hong Kong Polytechnic University, Hong Kong S.A.R.${}^{1}$
School of Cyber Science and Technology, Sun Yat-Sen University, Shenzhen, China.${}^{2}$
@@ 1.   Introduction
Utilizing sentence representations yielded by pre-trained language models (PLMs) to derive OOD confidence scores has been the de facto methodology for textual OOD detection. Specifically, PLMs are first fine-tuned on the ID data and then OOD detectors are applied on the sentence representations generated by PLMs. A body of works Hendrycks et al. (2020); Podolskiy et al. (2021); Uppaal et al. (2023) have shown that Transformer-based models can produce better sentence representations for OOD detection. However, these studies have mainly focused on evaluating the OOD detection performance of small-scale encoder-based Transformers, such as RoBERTa and BERT.
@@
Recently, large language models (LLMs) have made significant strides in various cognitive tasks, yet their capabilities on OOD detection remain largely unexplored. Unlike relatively small-scale PLMs used by prior studies, LLMs often display notable differences. In particular, the majority of previously prominent PLMs utilized for OOD detection adopt the encoder-based architecture, such as BERT and RoBERTa. These models are predominantly designed with a pre-training objective that focuses on sentence classification. However, recent LLMs Touvron et al. (2023); Zeng et al. (2022); Du et al. (2022); Chowdhery et al. (2022); Chung et al. (2022) exclusively adopt an autoregressive training objective during pre-training. Consequently, the hidden states of LLMs are specialized for next token prediction, which could influence their performance in OOD detection. Moreover, previous works test changes in OOD detection when adapting PLMs to downstream tasks through discriminative fine-tuning, even for decoder-based models Cho et al. (2023). However, a more intuitive approach is to probe the pre-training knowledge of LLMs through generative fine-tuning, which better aligns LLMs’ pre-training objective with downstream tasks. Thus, it is imperative to extensively investigate the OOD detection capabilities of LLMs to gain deeper insights into their potential and limitations.
@@ 2.1.   Out-of-Distribution Detection
Out-of-Distribution (OOD) detection has a long history in machine learning and is highly related to research topics like outlier detection, anomaly detection and novelty detection Hendrycks and Mazeika (2022). In the task setting of OOD detection, the in-distribution is characterized by the labeled training dataset and the out-of-distribution refers to anything else that possesses distributional shifts. OOD detection distinguishes from other related topics in the case that it requires both of the ID and OOD classification accuracy Yang et al. (2021).
@@ 2.2.   Textual OOD Detection with PLMs
Several works have further evaluated the OOD performance of PLMs with respect to commonly used OOD detectors including MSP Hendrycks and Gimpel (2017), Mahalanobis distance (MD) Lee et al. (2018), and Energy sore Liu et al. (2020). For example, Podolskiy et al. (2021) show that the Gaussian distribution assumption of MD matches better the representation space of BERT and can yield the best OOD performance in intent OOD detection benchmarks. Zhou et al. (2021) show that a contrastive regularizer can further improve the sentence representation of Transformers for OOD detection.
More recently, Uppaal et al. (2023) present a thorough analysis on the fine-tuning strategies for OOD detection with RoBERTa and show that RoBERTa Liu et al. (2019) without fine-tuning can achieve near-perfect OOD performance. Similarly, we find that LLMs can also achieve perfect OOD performance without fine-tuning. Moreover, using different fine-tuning objectives will have different effects on LLMs. Cho et al. (2023) explore the OOD detection capability of medium-sized PLMs (such as GPT-2 Radford et al. (2019)), as well as the impact of various ID fine-tuning techniques. While they also assess decoder-based models, the models they evaluated are not extensive as this work and they neglect to undertake generative ID tuning, a crucial step to fine-tune decoder-based models for downstream ID tasks. Furthermore, the models they examine remain at relatively moderate scales, and an exploration of the possible data-efficient characteristics intrinsic to the model is lacking in their study.
In recent months, large language models (LLMs) have been leading a paradigm shift in the NLP community Touvron et al. (2023); Wei et al. (2022); Zeng et al. (2022); Du et al. (2022); Chowdhery et al. (2022); Chung et al. (2022); Feng et al. (2023). Solving NLP tasks in a generative manner with LLMs has swept almost all ML tasks. These LLMs commonly adopt the decoder-based architecture and are trained with the autoregressive objective. In this paper, we focus on the OOD performance of open-source LLMs and anticipate our work can provide useful insights for OOD detection under this paradigm.
@@ 3.2.   OOD Detection with LLMs
The overview of our OOD detection framework is illustrated in Figure 2. Our primary focus is on decoder-like LLMs, such as LLaMA Touvron et al. (2023), as they have demonstrated excellent performance when their model size scales up OpenAI (2023); Brown et al. (2020). To obtain a comprehensive observation, we conduct OOD detection experiments on two different semantic distribution settings Ming et al. (2022); Lin et al. (2021): far-OOD and near-OOD (cf. Figure 1). Regarding OOD detection methods, we focus on the prevailing post-hoc paradigm Yang et al. (2021). In the following, we elaborate on how to integrate the post-hoc OOD detectors into decoder-like LLMs, which has never been addressed in existing literature.
@@
Logits-based OOD score functions operate on the final class-related logits. In generative classification, the generated class name is usually composed of several tokens, e.g., “positive” consists of “posi” and “tive”. Instead of calculating the probability (logits) for the complete ID class name, we simplify the process by considering the probability assigned to the first token of its class name. For instance, in a sentiment analysis task with classes like "positive" and "negative" as depicted in Figure 2, we only need to identify the probability corresponding to the tokens “posi” and “negative” respectively. Considering that different class names may have common prefixes, such as “positive” and “position”, we will rephrase the conflicting class names at the beginning of ID training, such as replacing “position” with “location”. In practice, we observe this re-translation has no impact on the ID task. Overall, there are mainly two logits-based functions:
@@
  • Energy score (Energy) Liu et al. (2020); LeCun et al. (2006) computes confidence score $\mathcal{S}(x)=\log\sum_{i}^{K}e^{(w^{T}\cdot z)_{i}}$ where $w^{T}$ is the weight of the language model head and $z$ is all word embeddings. Note that for both MSP and Energy, we only select the probability and logits corresponding to the first token of each class name, as mentioned before.
Distance-based OOD score functions apply to sentence representations. Prior studies using encoder-based PLMs treated the embeddings of special token <cls> as sentence representations. For LLMs, we employ the embeddings of the last token as the representation. There are mainly two functions considered for evaluation: Mahalanobis distance (Maha) Lee et al. (2018) and Cosine distance (Cosine) Zhou et al. (2021) [^1] We refer authors to original papers for more details..
@@ 5.2.   OOD Detection with Generatively Fine-tuned LLMs
Besides, in Table 4, we compare encoder-based and decoder-based Transformers and observe significant improvement on decoder-based LLMs.
@@ 7.   Bibliographical References
  • Touvron et al. (2023) Hugo Touvron, Thibaut Lavril, Gautier Izacard, Xavier Martinet, Marie-Anne Lachaux, Timothée Lacroix, Baptiste Rozière, Naman Goyal, Eric Hambro, Faisal Azhar, et al. 2023. Llama: Open and efficient foundation language models. arXiv preprint arXiv:2302.13971.

### v4

@@ Abstract
Out-of-distribution (OOD) detection plays a vital role in enhancing the reliability of machine learning models. As large language models (LLMs) become more prevalent, the applicability of prior research on OOD detection that utilized smaller-scale Transformers such as BERT, RoBERTa, and GPT-2 may be challenged, due to the significant differences in the scale of these models, their pre-training objectives, and the paradigms used for inference. This paper initiates a pioneering empirical investigation into the OOD detection capabilities of LLMs, focusing on the LLaMA series ranging from 7B to 65B in size. We thoroughly evaluate commonly used OOD detectors, examining their performance in both zero-grad and fine-tuning scenarios. Notably, we alter previous discriminative in-distribution fine-tuning into generative fine-tuning, aligning the pre-training objective of LLMs with downstream tasks. Our findings unveil that a simple cosine distance OOD detector demonstrates superior efficacy, outperforming other OOD detectors. We provide an intriguing explanation for this phenomenon by highlighting the isotropic nature of the embedding spaces of LLMs, which distinctly contrasts with the anisotropic property observed in smaller BERT family models. The new insight enhances our understanding of how LLMs detect OOD data, thereby enhancing their adaptability and reliability in dynamic environments. We have released the source code at https://github.com/Awenbocc/LLM-OOD for other researchers to reproduce our results.
@@
Bo Liu1∗thanks:   Equal contribution., Li-Ming Zhan1∗, Zexin Lu1∗, Yujie Feng1, Lei Xue2, Xiao-Ming Wu1†thanks:   Corresponding author.
Department of Computing, The Hong Kong Polytechnic University, Hong Kong S.A.R.1
School of Cyber Science and Technology, Sun Yat-Sen University, Shenzhen, China.2
@@ 1.   Introduction
Utilizing sentence representations yielded by pre-trained language models (PLMs) to derive OOD confidence scores has been the de facto method for textual OOD detection. Specifically, PLMs are first fine-tuned on the ID data and then OOD detectors are applied on the sentence representations generated by PLMs. Compared to ID data, there are two types of OOD instances: far-OOD where ID and OOD data come from different domains and near-OOD where ID and OOD data come from the same domain but with different classes, as shown in Figure 1. Typically, near-OOD samples are harder to recognize. A body of works Hendrycks et al. (2020); Podolskiy et al. (2021); Uppaal et al. (2023); Zhan et al. (2024) have shown that Transformer-based models can produce better sentence representations for OOD detection. However, these studies have mainly focused on evaluating the OOD detection performance of small-scale encoder-based Transformers, such as RoBERTa and BERT.
@@
Recently, large language models (LLMs) have made significant strides in various cognitive tasks, yet their capabilities on OOD detection remain largely unexplored. Unlike relatively small-scale PLMs used by prior studies, LLMs often display notable differences. In particular, the majority of previously prominent PLMs utilized for OOD detection adopt the encoder-based architecture, such as BERT and RoBERTa. These models are predominantly designed with a pre-training objective that focuses on sentence classification. However, recent LLMs Touvron et al. (2023a); Zeng et al. (2022); Du et al. (2022); Chowdhery et al. (2022); Chung et al. (2022) exclusively adopt an autoregressive training objective during pre-training. Consequently, the hidden states of LLMs are specialized for next token prediction, which could influence their performance in OOD detection. Moreover, previous works test changes in OOD detection when adapting PLMs to downstream tasks through discriminative fine-tuning, even for decoder-based models Cho et al. (2023). However, a more intuitive approach is to probe the pre-training knowledge of LLMs through generative fine-tuning, which better aligns LLMs’ pre-training objective with downstream tasks. Thus, it is imperative to extensively investigate the OOD detection capabilities of LLMs to gain deeper insights into their potential and limitations.
@@ 2.1.   Out-of-Distribution Detection
Out-of-Distribution (OOD) detection has a long history in machine learning and is highly related to research topics like outlier detection, anomaly detection and novelty detection Hendrycks and Mazeika (2022). In the task setting of OOD detection, the in-distribution is characterized by the labeled training dataset and the out-of-distribution refers to anything else that possesses distributional shifts. Note that OOD detection differs from outlier detection in that it requires accurate classification of both of ID and OOD data Yang et al. (2021).
@@ 2.2.   Textual OOD Detection with PLMs
Several works have further evaluated the OOD performance of PLMs with respect to commonly used OOD detectors including MSP Hendrycks and Gimpel (2017), Mahalanobis distance (MD) Lee et al. (2018), and Energy score Liu et al. (2020). For example, Podolskiy et al. (2021) show that the Gaussian distribution assumption of MD better matches the representation space of BERT and can yield the best OOD performance in intent OOD detection benchmarks. Zhou et al. (2021) show that a contrastive regularizer can further improve the sentence representation of Transformers for OOD detection.
More recently, Uppaal et al. (2023) present a thorough analysis on the fine-tuning strategies for OOD detection with RoBERTa and show that RoBERTa Liu et al. (2019) without fine-tuning can achieve near-perfect far-OOD detection performance. Similarly, we find that LLMs can also achieve perfect far-OOD detection performance without fine-tuning. Cho et al. (2023) explore the OOD detection capability of medium-sized PLMs (such as GPT-2 Radford et al. (2019)), as well as the impact of various ID fine-tuning techniques. While they also assess decoder-based models, the models they evaluated are not extensive as this work and they neglect to undertake generative ID tuning, a crucial step to fine-tune decoder-based models for downstream ID tasks. Furthermore, the models they examine remain at relatively moderate scales, and an exploration of the possible data-efficient characteristics of the model is lacking.
Recently, large language models (LLMs) have been leading a paradigm shift in the field of natural language processing (NLP) Touvron et al. (2023a); Wei et al. (2022); Zeng et al. (2022); Du et al. (2022); Chowdhery et al. (2022); Chung et al. (2022); Lu et al. (2021); Feng et al. (2023). The use of LLMs to solve NLP tasks in a generative way has become widespread. These LLMs commonly adopt the decoder-based architecture and are trained with the autoregressive objective. In this paper, we focus on the OOD performance of open-source LLMs and anticipate our work can provide useful insights for OOD detection under this paradigm.
@@ 3.2.   OOD Detection with LLMs
The overview of our OOD detection framework is illustrated in Figure 2. Our primary focus is on decoder-like LLMs, such as LLaMA Touvron et al. (2023a), as they have demonstrated excellent performance when their model size scales up OpenAI (2023); Brown et al. (2020). To obtain a comprehensive observation, we conduct OOD detection experiments on two different semantic distribution settings Ming et al. (2022); Lin et al. (2021): far-OOD and near-OOD (cf. Figure 1). Regarding OOD detection methods, we focus on the prevailing post-hoc paradigm Yang et al. (2021). In the following, we elaborate on how to integrate post-hoc OOD detectors into decoder-style LLMs, which has not been addressed in existing literature.
@@
Logits-based OOD score functions operate on the final class-related logits. In generative classification, the generated class name is usually composed of several tokens, e.g., “positive” consists of “posi” and “tive”. Instead of calculating the probability (logits) for the entire ID class name, we simplify the process by considering the probability assigned to the first token of its class name. For instance, in a sentiment analysis task with classes like “positive” and “negative” as depicted in Figure 2, we only need to identify the probability corresponding to the tokens “posi” and “negative” respectively. Considering that different class names may have common prefixes, such as “positive” and “position”, we will rephrase the conflicting class names at the beginning of ID training, such as replacing “position” with “location”. In practice, we observe this re-translation has no impact on the ID task. Overall, there are mainly two logits-based functions:
@@
  • Energy score (Energy) Liu et al. (2020); LeCun et al. (2006) computes confidence score $\mathcal{S}(x)=\log\sum_{i}^{K}e^{(w^{T}\cdot z)_{i}}$ where $w^{T}$ is the weight of the language model head and $z$ is all word embeddings. Note that for both MSP and Energy, we only select the probability and logits corresponding to the first token of each class name, as mentioned above.
Distance-based OOD score functions
apply to sentence representations. Prior studies using encoder-based PLMs treated the embeddings of special token <cls> as sentence representations. For LLMs, we employ the embeddings of the last token as the representation. There are mainly two functions considered for evaluation: Mahalanobis distance (Maha) Lee et al. (2018) and Cosine distance (Cosine) Zhou et al. (2021) [^1] We refer authors to original papers for more details..
@@ 5.2.   OOD Detection with Generatively Fine-tuned LLMs
Besides, in Table 4, we compare encoder-based and decoder-based Transformers and observe impressive improvement on decoder-based LLMs.
@@
6.   Analysis
6.1.   Performance of More LLMs
ID Dataset PTMs Maha Cosine MSP Energy
Zero-grad
SST-2 OPT-6.7B 0.982 0.983 0.413 0.571
LLaMA-7B 0.991 0.990 0.905 0.368
LLaMA2-7B 0.991 0.997 0.917 0.516
Fine-tuned
CLINC-Banking (Full) OPT-6.7B 0.921 0.932 0.915 0.922
LLaMA-7B 0.958 0.964 0.936 0.930
LLaMA2-7B 0.967 0.970 0.944 0.937
Table 6: OOD detection performance of various LLMs. AUROC result is reported.
Additionally, we test OPT-6.7B Zhang et al. (2022) and LLaMA2-7B Touvron et al. (2023b) here with settings that zero-grad OOD performance for SST-2 (ID) and fine-tuned OOD performance for CLINC-Banking, as shown in Table 6. For simplicity, we report the AUROC result of each OOD score function. Overall, the OOD detection performance of LLaMA2-7B surpasses that of LLaMA-7B, and LLaMA-7B outperforms OPT-6.7B, which is consistent with the general performance trends observed in these models.
6.2.   Impact of Quantization
SST-2
Quantized Maha Cosine MSP Energy
float32 0.991 0.990 0.905 0.368
float16 0.990 0.987 0.893 0.331
Int8 0.955 0.960 0.874 0.228
Table 7: Impact of quantization for OOD detection.
We test the zero-grad OOD detection performance of LLaMA-7B with different quantized levels (float32, float16, and Int8) for SST-2 as the ID task, shown in Table 7. It was observed that the float16 model can largely preserve the OOD detection capability of the model. However, 8-bit quantization leads to a degradation in its capability.
6.3.   Error Analysis
We mainly analyze the fine-tuning OOD detection performance in the near-OOD setting since both RoBERTa Liu et al. (2019) and LLaMA Touvron et al. (2023a) can achieve near-perfect performance in the far-OOD setting. Here, we provide the OOD detection performance with fine-tuned RoBERTa in the following Table 8, as a complement to Table 3.
CLINC-Banking (Full)
PTMs Maha Cosine MSP Energy
LLaMA-7B 0.958 0.964 0.936 0.930
RoBERTa-L 0.821 0.793 0.670 0.717
Table 8: Performance comparison between LLaMA and RoBERTa-L in the near-OOD setting. AUROC result is reported for each score function.
Overall, LLaMA’s detection capabilities are much stronger than RoBERTa’s. Through the analysis of specific error cases, we found that RoBERTa struggles to distinguish semantically similar OOD samples. For instance, when choosing the “bank_balance” class (e.g., the sentence “what’s my account balance”) as the ID distribution, RoBERTa tends to incorrectly classify the majority of “bill_balance” class inputs (such as “what are my bills this month”) as “bank_balance”. In contrast, LLaMA generally makes correct judgments in most cases. We attribute this phenomenon to the anisotropy characteristic which is explained in Section 5.2), i.e., sentence embeddings produced by the BERT family models have been noted to possess an undesirable characteristic of being concentrated within a narrow cone, causing representation degeneration. Despite this, in the cases involving extremely similar semantics, LLaMA also makes errors in judgment, such as misclassifying “bill_balance: what is the balance on my bills” as “bank_balance”.
@@ 7.   Conclusion
Limitations and Future Work: In this work, we only tested textual LLMs and ignored the exploration for multi-modal LLMs Liu et al. (2024), which can be seen as a limitation. More OOD detection in multi-modal scenarios (including both general and medical domains) Liu et al. (2021); Radford et al. (2021); Zhu et al. (2023); Liu et al. (2022); Zhang et al. (2023) will be conducted in the future.
Acknowledgments
We thank the anonymous reviewers for their valuable feedback. This research was partially supported by the grant of HK ITF ITS/359/21FP.
@@ 8.   Bibliographical References
  • Liu et al. (2021) Bo Liu, Li-Ming Zhan, Li Xu, Lin Ma, Yan Yang, and Xiao-Ming Wu. 2021. Slake: A semantically-labeled knowledge-enhanced dataset for medical visual question answering. In 2021 IEEE 18th International Symposium on Biomedical Imaging (ISBI), pages 1650–1654. IEEE.
  • Liu et al. (2022) Bo Liu, Li-Ming Zhan, Li Xu, and Xiao-Ming Wu. 2022. Medical visual question answering via conditional reasoning and contrastive learning. IEEE transactions on medical imaging, 42(5):1532–1545.
  • Liu et al. (2024) Haotian Liu, Chunyuan Li, Qingyang Wu, and Yong Jae Lee. 2024. Visual instruction tuning. Advances in neural information processing systems, 36.
@@
  • Lu et al. (2021) Zexin Lu, Jing Li, Yingyi Zhang, and Haisong Zhang. 2021. Getting your conversation on track: Estimation of residual life for conversations. In 2021 IEEE Spoken Language Technology Workshop (SLT), pages 1036–1043. IEEE.
@@
  • Radford et al. (2021) Alec Radford, Jong Wook Kim, Chris Hallacy, Aditya Ramesh, Gabriel Goh, Sandhini Agarwal, Girish Sastry, Amanda Askell, Pamela Mishkin, Jack Clark, et al. 2021. Learning transferable visual models from natural language supervision. In International conference on machine learning, pages 8748–8763. PMLR.
@@
  • Touvron et al. (2023a) Hugo Touvron, Thibaut Lavril, Gautier Izacard, Xavier Martinet, Marie-Anne Lachaux, Timothée Lacroix, Baptiste Rozière, Naman Goyal, Eric Hambro, Faisal Azhar, et al. 2023a. Llama: Open and efficient foundation language models. arXiv preprint arXiv:2302.13971.
  • Touvron et al. (2023b) Hugo Touvron, Louis Martin, Kevin Stone, Peter Albert, Amjad Almahairi, Yasmine Babaei, Nikolay Bashlykov, Soumya Batra, Prajjwal Bhargava, Shruti Bhosale, et al. 2023b. Llama 2: Open foundation and fine-tuned chat models. arXiv preprint arXiv:2307.09288.
@@
  • Zhan et al. (2024) Li-Ming Zhan, Bo Liu, and Xiao-Ming Wu. 2024. Vi-ood: A unified representation learning framework for textual out-of-distribution detection. arXiv preprint arXiv:2404.06217.
  • Zhang et al. (2023) Hao Zhang, Hongyang Li, Feng Li, Tianhe Ren, Xueyan Zou, Shilong Liu, Shijia Huang, Jianfeng Gao, Lei Zhang, Chunyuan Li, et al. 2023. Llava-grounding: Grounded visual chat with large multimodal models. arXiv preprint arXiv:2312.02949.
  • Zhang et al. (2022) Susan Zhang, Stephen Roller, Naman Goyal, Mikel Artetxe, Moya Chen, Shuohui Chen, Christopher Dewan, Mona Diab, Xian Li, Xi Victoria Lin, et al. 2022. Opt: Open pre-trained transformer language models. arXiv preprint arXiv:2205.01068.
@@
  • Zhu et al. (2023) Deyao Zhu, Jun Chen, Xiaoqian Shen, Xiang Li, and Mohamed Elhoseiny. 2023. Minigpt-4: Enhancing vision-language understanding with advanced large language models. arXiv preprint arXiv:2304.10592.